# Uncertainty quantification for data-driven weather models

### v1

@@ 4.2.2 Distributional regression network (DRN)
A key limitation of the EasyUQ approach is that there is no straightforward way to include additional predictor variables besides the deterministic forecasts of the target variable of interest. However, recent research on ML-based ensemble post-processing methods has highlighted that incorporating additional predictors is a key aspect in the substantial improvements achieved by these approaches (Rasp and Lerch, 2018). Therefore, we consider a DRN, a parametric neural network (NN)-based approach first proposed in Rasp and Lerch (2018) as an alternative PH method. To introduce DRN, we slightly extend the notation from above and use $\mathbb{X}_{l,t,h}^{\text{PW}}$ to denote the (vector of) deterministic Pangu-Weather forecasts for all considered output variables at location $l\in\mathbb{L}_{E}$, initialization time $t$ and lead time $h$. Note that we consider the output variables from the corresponding vertical level only, and leave extensions towards incorporating additional inputs from other levels for future work.
Based on the deterministic model output, $\mathbb{X}_{l,t,h}^{\text{PW}}$, the DRN approach proceeds by training a NN which yields the parameters of a suitable parametric distribution for the target variable as its output. DRN enables the use of arbitrary predictors as inputs to the NN, including additional meteorological variables (in our case from Pangu-Weather outputs) and location information. The NN parameters are determined by minimizing the CRPS over the training dataset. As for EasyUQ, we estimate separate models for every lead time, but note that considering multiple lead times jointly can be a viable alternative (Primo et al., 2024). In our case study, we use a Gaussian predictive distribution for all target variables and closely follow Rasp and Lerch (2018) in our implementation. We fit a single DRN model for each forecast horizon $h$ jointly over all grid points $l\in\mathbb{L}_{E}$ in the target domain. In addition to the deterministic Pangu-Weather forecasts $\mathbb{X}_{l,t,h}^{\text{PW}}$, the locations are encoded via a positional embedding that maps a location $l\in\mathbb{L}_{E}$ to a vector of latent features, which are then used as auxiliary input variables of the NN. This procedure aims at making the model locally adaptive, while avoiding the training of a separate model at every grid point. The DRN model thus yields a predictive distribution
@@ 4.3 Forecast evaluation methods
where $\mathbbm{1}$ denotes the indicator function, and $F$ is assumed to have a finite first moment. Closed-form expressions of the CRPS are available for both sample-based predictive distributions in the form of an $M$-member ensemble, as well as for many parametric families of forecast distributions, see, e.g., Jordan et al. (2019).
@@ 5 Results
Table 1: Mean CRPS of all methods and variables across the spatial domain for three different groups of lead times, with the best-performing method highlighted in bold. Note that the CRPS values for Z500 are scaled by a factor of $0.01$.
@@
We here compare the previously introduced UQ methods based on their out-of-sample predictive performance. The ensemble forecasts from the operational 50-member ECMWF model are used as a baseline, and can be considered as a state-of-the-art physics-based NWP ensemble model. Note that we did not apply any post-processing to the ECMWF ensemble forecasts here, see the discussion in Section 6 for details. The evaluation and training setup follows the descriptions in Section 2, and we generate ensembles of size $M=50$ for all initial condition ensemble approaches. The PH methods are evaluated based on their predictive distributions, i.e., the empirical CDF for EasyUQ and the Gaussian forecast distribution for DRN.
@@
Table 1 provides the mean CRPS for all UQ methods, averaged over all grid points and stratified into three groups of forecast lead times. Figure 3 shows the mean CRPS as a function of the forecast lead time for all variables. Both illustrations indicate that the use of data-driven weather forecasts in concert with the PH methods proposed here can yield improvements over the ECMWF ensemble forecasts. The extent of these improvements, and the relative performance of the different UQ methods, strongly depends on the variable of interest as well as the forecast lead time. Generally, DRN yields the best forecasts at shorter lead times, followed closely by the EasyUQ model. For longer lead times up to 120 h, the CRPS of the ECMWF ensemble is similar to that of the DRN, EasyUQ and RFP approaches, and for lead times beyond 120 h, the ECMWF ensemble performs better than all compared UQ methods. The most pronounced differences and most clear improvements over the ECMWF ensemble forecasts can be observed for T2M. The rankings among the different UQ methods are mostly identical across the considered target variables, with the PH methods (DRN and EasyUQ) showing better forecasts at shorter lead times, whereas the RFP approach yields the best forecasts at longer lead times. Interestingly and in contrast to previous studies on ensemble post-processing (e.g., Schulz and Lerch, 2022), DRN only yields relatively minor improvements over the considerably simpler EasyUQ method. A potential explanation might be that DRN here is restricted to much fewer additional predictor variables compared to other studies. Therefore, further improvements might be achieved by, e.g., incorporating Pangu-Weather outputs from other vertical levels. The GNP and IFSP approaches lead to substantially worse forecasts compared to the other methods for all variables, in particular for z500.
@@
Figure 4 shows the CRPSS of the different UQ methods over the spatial domain, using the ECMWF ensemble as a reference forecast. For most methods, target variables and lead times, there are some geographical regions where improvements over the ECMWF ensemble are obtained. The most notable improvements and variations can be observed for the variable T2M. There, the improvements over the ECMWF ensemble forecasts are most pronounced over land, and even for a lead time of 168 h, all methods show a positive skill score over mountainous regions. In particular, the PH methods indicate a notably better performance. A generally similar spatial pattern, albeit with less pronounced improvements over the ECMWF ensemble, can be observed for U10. The areas with positive CRPSS values for Z500 seem to be more concentrated around the Mediterranean and south-eastern Europe, and the GNP and IFSP methods show notable worse performance.
@@
A complementary perspective on calibration is provided by the spread-skill plots in Figure 6, which shows the relationship between the RMSE of the mean forecast and the standard deviation of the ensemble predictions of the different UQ methods for the two temperature variables. For a well-calibrated ensemble forecast, the temporal evolution of the RMSE and the ensemble spread should roughly be equal (Fortin et al., 2014). As Figure 6 indicates, this is not the case for most of the UQ methods considered here  [^6] In order to make the post-processing methods comparable, the predicted mean and standard deviation at each gridpoint was extracted from the forecasts to compute the RMSE and spread, respectively.. In particular for the IC methods, the standard deviation is notably lower than the RMSE, indicating a clear underestimation of the true forecast uncertainty by these approaches. It is noteworthy that the standard deviation of the IFSP forecasts actually decreases during the first 24 h. This is somewhat surprising as the singular vectors at ECMWF represent the fastest growing perturbations over an optimization time window of 48 hours. Thus, a growth of the standard deviation would be expected as it is also the case in IFSP forecasts. A similar behavior with initially slow perturbation growth was already documented in Selz and Craig (2023) when Pangu-Weather was initialized from the members from the ECMWF ensemble data assimilation. We attribute the different standard deviation growth between RFP and IFSP to the fact that the RFP method leads to perturbations which are larger in scale than the IFSP perturbations (Fig. 2) and thus grow faster initially. The two post-hoc methods show a similar behavior, with a slight underdispersion but in general good calibration, for short lead times even better than the ECMWF ensemble forecast. It should be kept in mind though that the ECMWF ensemble forecasts also contain stochastic perturbed parametrization tendencies, which lead to a better calibrated ensemble.
@@
Figure 7 shows the mean bias for forecasts of selected target variables for all UQ methods. A strong negative bias in the IFSP forecasts is apparent already at shorter lead times for T2M and, in particular, Z500. The likely cause of this bias, which, in turn, explains the observed bad performance of the IFSP approach, are systematic differences between the initial conditions of the operational ECMWF ensemble and the ERA5 reanalysis fields, which we considered as ground truth. Further, there likely are differences in the representation of topography in the operational version of the IFS model and ERA5 due to differences in the underlying grid spacing. One pathway to addressing this, which has been suggested for example in Rasp et al. (2023), might be to evaluate the NWP-based forecasts (including, potentially, the IFSP forecasts) against the operational analysis. Although both post-processing methods operate separately on every grid point, only the DRN approach shows a strong granular pattern, while the bias pattern of the EasyUQ method appears notably more smooth and comparable to that of the RFP approach. The GNP methods shows fairly small biases which are on a comparable level to those of the better-performing RFP and DRN approaches. While no substantial differences in the overall level of the bias among these methods can be observed, the DRN forecasts for Z500 at a lead time of 168 hours interestingly shows the most pronounced biases among all compared methods.
@@ 6 Discussion and conclusions
To the best of our knowledge, our study is the first systematic comparison of different UQ methods to generate probabilistic weather forecasts from the deterministic data-driven weather model Pangu-Weather (Bi et al., 2023). The UQ approaches can be divided into IC-based methods, where an ensemble forecast is generated by initializing Pangu-Weather model runs from different sets of initial conditions, and PH methods, which operate on deterministic Pangu-Weather forecasts and generate probabilistic forecasts from the deterministic model inputs, based on past forecasts and corresponding observations. Overall, our results suggest that most of the UQ methods are able to provide probabilistic forecasts competitive with the operational (non-post-processed) ECMWF ensemble forecast.
@@
The IC approaches have the advantage that they generate realistic spatial forecast fields, as the input is perturbed over the whole spatial domain, whereas the PH methods generate a separate predictive distribution at every grid point. In particular, the RFP approach seems promising in that IC ensembles can be straightforwardly generated from past observation data with minimal tuning. However, the use of the IC methods comes at the cost of having to run the deterministic data-driven weather model multiple times, which can be demanding in terms of the computing and disk space requirements, in particular for generating and storing global high-resolution ensemble forecasts. By contrast, PH methods require a training dataset of past forecasts from the deterministic data-driven weather model and corresponding observations. While they have the advantage of potentially correcting systematic errors such as biases of the underlying deterministic model, additional modeling steps are required to generate spatially coherent forecast fields. A variety of two-step methods for multivariate post-processing is available, where in a first step, forecasts are post-processed separately at every grid point or lead time (using methods like, e.g., EasyUQ or DRN). In a second step, multivariate (e.g., spatial or temporal) dependencies are introduced by re-ordering samples from the univariate forecast distribution according to a dependence template via the use of copula functions. Popular approaches include the use of empirical copulas based on the physics-based NWP ensemble models (ensemble copula coupling, ECC; Schefzik et al., 2013), or based on past observations (Schaake shuffle; Clark et al., 2004). Comprehensive comparisons are for example available in Lerch et al. (2020) and Lakatos et al. (2023). In the context of post-processing data-driven weather model forecasts, the use of the ECC method comes with the benefit of obtaining a hybrid combination of a data-driven model producing the univariate forecasts at each grid point, and a physics-based ensemble model that provides information on the spatio-temporal dependencies. Recently proposed multivariate post-processing methods based on generative ML (Chen et al., 2024) further have the potential to better utilize various sources of input information and improve the multivariate probabilistic forecasts.
@@
The main objective of our study was to provide a general proof of concept for how to generate probabilistic forecasts from deterministic data-driven weather models. It should be seen as a first step into the direction of probabilistic data-driven weather models, and our results provide several avenues for further generalization and analysis. A natural benchmark method, which is missing from our comparisons, is the combination of the ECMWF ensemble forecasts with post-processing. Correcting systematic errors in NWP ensemble predictions was the original motivation for the development of post-processing methods such as DRN, and notable improvements in terms of the CRPS have been observed in numerous studies (Vannitsem et al., 2021). We were not able to include post-processed ECMWF ensemble forecasts since it proved difficult to obtain a suitably large archive of past forecasts due to exceedingly slow download processes. Nonetheless, this comparison would be of substantial interest. Recently, Bremnes et al. (2023) compare post-processed Pangu-Weather and physics-based weather forecasts on a station dataset over Norway and find that the forecast quality tends to be very similar after post-processing.
@@
Further, our comparisons have been focused on selected target variables, on using the gridded ERA5 data as ground truth, and on the CRPS as main evaluation metric. Operational weather services tend to evaluate their forecasts against the model’s own operational analysis, station observations (Rasp et al., 2023), and comprehensive comparisons of the UQ methods constitute an interesting starting point for future research. Over Europe, suitable station observation data has for example been collected within the EUPPBench benchmark dataset for post-processing (Demaeyer et al., 2023). Another important open question regarding the potential and limitations of data-driven weather models is whether they can reliably predict extreme weather events. Therefore, a targeted evaluation of the UQ methods in this regard, e.g., using proper weighted scoring rules (Lerch et al., 2017), represents another important direction for future model comparisons.
@@
The large data volumes and high dimensionality of global gridded predictions further poses a challenge regarding the scalability of ML-based post-processing methods such as DRN, for which it is an open question whether they will generalize well to global high-resolution forecasts. This calls for the development of new spatial post-processing methods which operate on the spatial forecast fields directly and are able to leverage predictive information present in the spatial structures. Over the past years, several approaches have been proposed, which utilize convolutional neural network or transformer architectures to enable probabilistic post-processing of spatial forecast fields (e.g., Grönquist et al., 2021; Ashkboos et al., 2022; Chapman et al., 2022; Ben-Bouallègue et al., 2023b; Horat and Lerch, 2024). The recently introduced WeatherBench 2 dataset (Rasp et al., 2023) provides a useful framework for comparisons.

### v2

@@ 4.2.2 Distributional regression network (DRN)
A key limitation of the EasyUQ approach is that there is no straightforward way to include additional predictor variables besides the deterministic forecasts of the target variable of interest. However, recent research on ML-based ensemble post-processing methods has highlighted that incorporating additional predictors is a key aspect in the substantial improvements achieved by these approaches (Rasp and Lerch, 2018). Therefore, we consider a DRN, a parametric neural network (NN)-based approach first proposed in Rasp and Lerch (2018) as an alternative PH method. To introduce DRN, we slightly extend the notation from above and use $\mathbb{X}_{l,t,h}^{\text{PW}}$ to denote the (vector of) deterministic Pangu-Weather forecasts for all considered output variables at location $l\in\mathbb{L}_{E}$, initialization time $t$ and lead time $h$. Note that we consider only predictor variables in the DRN from the same vertical atmospheric level as the target variable of interest. Extensions towards incorporating additional predictors from other vertical levels is left for future work.
Based on the deterministic model output, $\mathbb{X}_{l,t,h}^{\text{PW}}$, the DRN approach proceeds by training a NN which yields the parameters of a suitable parametric distribution for the target variable as its output. DRN enables the use of arbitrary predictors as inputs to the NN, including additional meteorological variables (in our case from Pangu-Weather outputs) and location information (i.e., latitude and longitude). The NN parameters are determined by minimizing the CRPS over the training dataset. As for EasyUQ, we estimate separate models for every lead time, but note that considering multiple lead times jointly can be a viable alternative (Primo et al., 2024). In our case study, we use a Gaussian predictive distribution for all target variables and closely follow Rasp and Lerch (2018) in our implementation. We fit a single DRN model for each forecast horizon $h$ jointly over all grid points $l\in\mathbb{L}_{E}$ in the target domain. In addition to the deterministic Pangu-Weather forecasts $\mathbb{X}_{l,t,h}^{\text{PW}}$, the locations are encoded via a positional embedding that maps a location $l\in\mathbb{L}_{E}$ to a vector of latent features, which are then used as auxiliary input variables of the NN. This procedure aims at making the model locally adaptive, while avoiding the training of a separate model at every grid point. The DRN model thus yields a predictive distribution
@@ 4.3 Forecast evaluation methods
where $\mathbb{I}$ denotes the indicator function, and $F$ is assumed to have a finite first moment. Closed-form expressions of the CRPS are available for both sample-based predictive distributions in the form of an $M$-member ensemble, as well as for many parametric families of forecast distributions, see, e.g., Jordan et al. (2019).
@@ 5 Results
5.1 UQ methods applied to Pangu-Weather
Table 1: Mean CRPS of all methods and variables across the European domain for three different groups of lead times, with the best-performing method highlighted in bold. Note that the CRPS values for Z500 are scaled by a factor of $0.01$.
@@ 5.1 UQ methods applied to Pangu-Weather
We here compare the previously introduced UQ methods based on their out-of-sample predictive performance. The ensemble forecasts from the operational 50-member ECMWF model are used as a baseline, and can be considered as a state-of-the-art physics-based NWP ensemble model. Note that we did not apply any post-processing to the ECMWF ensemble forecasts here, but compare selected UQ methods to post-processed deterministic IFS forecasts in the next subsection. The evaluation and training setup follows the descriptions in Section 2, and we generate ensembles of size $M=50$ for all initial condition ensemble approaches. The PH methods are evaluated based on their predictive distributions, i.e., the empirical CDF for EasyUQ and the Gaussian forecast distribution for DRN.
@@
Table 1 provides the mean CRPS for all UQ methods, averaged over all grid points in the European domain and stratified into three groups of forecast lead times. Figure 3 shows the mean CRPS as a function of the forecast lead time for all variables. Both illustrations indicate that the use of data-driven weather forecasts in concert with the PH methods proposed here can yield improvements over the ECMWF ensemble forecasts. The extent of these improvements, and the relative performance of the different UQ methods, strongly depends on the variable of interest as well as the forecast lead time. Generally, DRN yields the best forecasts at shorter lead times, followed closely by the EasyUQ model. For longer lead times up to 120 h, the CRPS of the ECMWF ensemble is similar to that of the DRN, EasyUQ and RFP approaches, and for lead times beyond 120 h, the ECMWF ensemble performs better than all compared UQ methods. The most pronounced differences and most clear improvements over the ECMWF ensemble forecasts can be observed for T2M. The rankings among the different UQ methods are mostly identical across the considered target variables, with the PH methods (DRN and EasyUQ) showing better forecasts at shorter lead times, whereas the RFP approach yields the best forecasts at longer lead times. Interestingly and in contrast to previous studies on ensemble post-processing (e.g., Schulz and Lerch, 2022), DRN only yields relatively minor improvements over the considerably simpler EasyUQ method. A potential explanation might be that DRN here is restricted to much fewer additional predictor variables compared to other studies. Therefore, further improvements might be achieved by, e.g., incorporating Pangu-Weather outputs from vertical atmospheric levels other than the vertical level of the target variable itself. The GNP and IFSP approaches lead to substantially worse forecasts compared to the other methods for all variables, in particular for Z500.
@@
Figure 4 shows the CRPSS of the different UQ methods over the spatial domain, using the ECMWF ensemble as a reference forecast. For most methods, target variables and lead times, there are some geographical regions where improvements over the ECMWF ensemble are obtained. The most notable improvements and variations can be observed for the variable T2M. There, the improvements over the ECMWF ensemble forecasts are most pronounced over land, and even for a lead time of 168 h, all methods show a positive skill score over mountainous regions. In particular, the PH methods indicate a notably better performance. A generally similar spatial pattern, albeit with less pronounced improvements over the ECMWF ensemble, can be observed for U10. The areas with positive CRPSS values for Z500 seem to be more concentrated around the Mediterranean and south-eastern Europe, and the GNP and IFSP methods perform notably worse.
@@
A complementary perspective on calibration is provided by the spread-skill plots in Figure 6, which shows the relationship between the RMSE of the mean forecast and the standard deviation of the ensemble predictions of the different UQ methods for the two temperature variables. For a well-calibrated ensemble forecast, the average ensemble spread should be roughly equal to the RMSE of the ensemble mean predictions for each lead time (Fortin et al., 2014). As Figure 6 indicates, this is not the case for most of the UQ methods considered here  [^5] In order to make the post-processing methods comparable, the predicted mean and standard deviation at each grid point were extracted from the forecasts to compute the RMSE and spread, respectively.. In particular for the IC methods, the standard deviation is notably lower than the RMSE, indicating a clear underestimation of the true forecast uncertainty by these approaches. It is noteworthy that the standard deviation of the IFSP forecasts actually decreases during the first 24 h. This is somewhat surprising as the singular vectors at ECMWF represent the fastest growing perturbations over an optimization time window of 48 hours. Thus, a growth of the standard deviation would be expected. A similar behavior with initially slow perturbation growth was already documented in Selz and Craig (2023) when Pangu-Weather was initialized with rescaled perturbations from the members of the ECMWF ensemble data assimilation. We attribute the different standard deviation growth between RFP and IFSP to the fact that the RFP method leads to perturbations which are larger in scale and magnitude than the IFSP perturbations (Fig. 2) and thus grow faster initially. The two post-hoc methods show a similar behavior, with a slight underdispersion but in general good calibration, for short lead times even better than the ECMWF ensemble forecast. It should be kept in mind though that the ECMWF ensemble forecasts also contain the stochastically perturbed parametrization tendency scheme (Buizza1999), which leads to a better calibrated ensemble.
@@
Figure 7 shows the mean bias for forecasts of selected target variables for all UQ methods, where the bias is computed as the difference between the realizing observation and the mean forecast. A strong negative bias in the IFSP forecasts is apparent already at shorter lead times for T2M and, in particular, Z500. The likely cause of this bias, which, in turn, explains the observed bad performance of the IFSP approach, are systematic differences between the initial conditions of the operational ECMWF ensemble and the ERA5 reanalysis fields, which we considered as ground truth. Further, there are differences in the representation of topography in the operational version of the IFS model and ERA5 due to differences in the underlying grid spacing. One future pathway to addressing this, which has been suggested for example in Rasp et al. (2024), is to evaluate the NWP-based forecasts (including, potentially, the IFSP forecasts) against the operational analysis. Although both post-processing methods operate separately on every grid point, only the DRN approach shows a strong granular pattern, while the bias pattern of the EasyUQ method appears notably more smooth and comparable to that of the RFP approach. The GNP method shows fairly small biases which are on a comparable level to those of the better-performing RFP and DRN approaches. While no substantial differences in the overall level of the bias among these methods can be observed, the DRN forecasts for Z500 at a lead time of 168 hours interestingly show the most pronounced biases among all compared methods.
5.2 Comparison to post-processed IFS forecasts
Table 2: Mean CRPS for the post-processing methods applied to IFS and Pangu-Weather predictions across the spatial domain for three different groups of lead times, with the best-performing method highlighted in bold. Note that the CRPS values for Z500 are scaled by a factor of $0.01$.
Variable IFS IFS+EUQ IFS+DRN Pangu+EUQ Pangu+DRN
6h - 48h U10 0.54 0.53 0.51 0.53 0.51
V10 0.54 0.53 0.51 0.53 0.51
T2M 0.57 0.47 0.44 0.60 0.41
T850 0.43 0.48 0.42 0.57 0.41
Z500 0.33 0.33 0.30 0.36 0.32
48h - 120h U10 0.96 0.98 0.96 1.05 1.03
V10 0.96 0.99 0.96 1.05 1.03
T2M 0.75 0.66 0.64 0.69 0.67
T850 0.75 0.78 0.75 0.82 0.79
Z500 1.21 1.27 1.22 1.35 1.29
$\geq$ 120h U10 1.54 1.57 1.55 1.70 1.68
V10 1.58 1.61 1.59 1.74 1.71
T2M 1.05 1.00 0.98 1.13 1.10
T850 1.33 1.38 1.35 1.55 1.48
Z500 2.91 3.03 2.97 3.36 3.25
Figure 8: CRPSS of the different T2M post-processing methods applied to IFS and Pangu-Weather predictions over the spatial domain, using the ECMWF ensemble as a reference method. The rows correspond to specific forecasting lead times. Note that positive CRPSS values indicate an improvement over the reference in terms of the CRPS at the respective grid point.
Correcting systematic errors in NWP ensemble predictions was the original motivation for the development of post-processing methods such as DRN, and notable improvements in terms of the CRPS have been observed in numerous studies (Vannitsem et al., 2021). As post-processing approaches have been widely adopted to improve physics-based NWP forecasts, post-processed IFS predictions thus constitute a natural benchmark for the UQ methods. To compare post-processed data-driven and physics-based weather models, we applied the EasyUQ and DRN approach to the deterministic IFS forecast obtained from WeatherBench 2 (Rasp et al., 2024), with the same specification and experimental setup as for the Pangu-Weather predictions. To enable a direct and fair comparison, we here utilize the deterministic IFS forecast only, and leave a comparison to post-processed ECMWF ensemble forecasts for future work.
Table 2 summarizes the mean CRPS values for a direct comparison of the post-processed IFS and Pangu-Weather predictions. For shorter lead times of 6–48 hours, applying DRN to Pangu-Weather shows minimally better results than DRN applied to the IFS predictions. More notable differences (in favor of the post-processed IFS predictions) can be observed for EasyUQ. For lead times between 48 and 120 hours, IFS+DRN achieves the best performance and post-processed IFS forecasts show clear improvements over post-processed Pangu-Weather forecasts. That said, the post-processed IFS predictions only offer clear improvements over the raw ECMWF ensemble forecasts for T2M. For lead times above 120 hours, the ECMWF ensemble forecasts perform best for most variables, with the exception of T2M where post-processing leads to improvements.
Figure 8 shows the CRPS skill score for the aforementioned post-processing methods for T2M. While post-processed Pangu-Weather forecasts obtain sizable improvements in high altitude areas and seem to perform better than post-processed IFS forecasts over land grid points for shorter lead times, their predictive performance over the sea is notably worse. This comparison highlights that skillful probabilistic forecasts can be obtained by applying post-processing to both physics-based and data-driven weather models, with the best-performing combination varying across the meteorological variable, lead time and location.
@@ 6 Discussion and conclusions
To the best of our knowledge, our study is the first systematic comparison of different UQ methods to generate probabilistic weather forecasts from the deterministic data-driven weather model Pangu-Weather (Bi et al., 2023). The UQ approaches can be divided into initial condition-based methods, where an ensemble forecast is generated by initializing Pangu-Weather model runs from different sets of initial conditions, and post-hoc methods, which operate on deterministic Pangu-Weather forecasts and generate probabilistic forecasts from the deterministic model inputs, based on past forecasts and corresponding observations. Overall, our results suggest that most of the UQ methods are able to provide probabilistic forecasts that are competitive with the operational (raw and post-processed) ECMWF ensemble forecast.
@@
Our evaluation has been restricted to separately considering individual meteorological variables and grid points, and does not take into account spatial or inter-variable dependencies. The IC approaches have the advantage that they generate realistic spatial forecast fields, as the input is perturbed over the whole spatial domain, whereas the PH methods generate a separate predictive distribution at every grid point. In particular, the RFP approach seems promising in that IC ensembles can be straightforwardly generated from past observation data with minimal tuning. However, the use of the IC methods comes at the cost of having to run the deterministic data-driven weather model multiple times, which can be demanding in terms of the computing and disk space requirements, in particular for generating and storing global high-resolution ensemble forecasts. By contrast, PH methods require a training dataset of past forecasts from the deterministic data-driven weather model and corresponding observations. While they have the advantage of potentially correcting systematic errors such as biases of the underlying deterministic model, additional modeling steps are required to generate spatially coherent forecast fields. A variety of two-step methods for multivariate post-processing is available, where in a first step, forecasts are post-processed separately at every grid point or lead time (using methods like, e.g., EasyUQ or DRN). In a second step, multivariate (e.g., spatial or temporal) dependencies are introduced by re-ordering samples from the univariate forecast distribution according to a dependence template via the use of copula functions. Popular approaches include the use of empirical copulas based on the physics-based NWP ensemble models (ensemble copula coupling, ECC; Schefzik et al., 2013), or based on past observations (Schaake shuffle; Clark et al., 2004). Comprehensive comparisons are for example available in Lerch et al. (2020) and Lakatos et al. (2023). In the context of post-processing data-driven weather model forecasts, the use of the ECC method comes with the benefit of obtaining a hybrid combination of a data-driven model producing the univariate forecasts at each grid point, and a physics-based ensemble model that provides information on the spatio-temporal dependencies. However, ECC would still require physics-based NWP ensemble forecasts, in contrast to using Schaake shuffle to determine decencies from past observations. Recently proposed multivariate post-processing methods based on generative ML (Chen et al., 2024) further have the potential to better utilize various sources of input information and improve the multivariate probabilistic forecasts.
@@
The main objective of our study was to provide a general proof of concept for how to generate probabilistic forecasts from deterministic data-driven weather models. It should be seen as a first step into the direction of probabilistic data-driven weather models, and our results provide several avenues for further generalization and analysis. As a natural benchmark, we also applied the PH methods to the physics-based IFS forecast. Our results indicate that at least for shorter lead times, the performance of post-processed IFS forecasts is in general quite similar to the post-processed data-driven forecast or sometimes even worse. These findings are in line with results of Bremnes et al. (2024), who compare post-processed Pangu-Weather and physics-based weather forecasts on a station dataset over Norway and find that the forecast quality tends to be very similar after post-processing. Comprehensive comparisons of post-processed data-driven and physics-based weather forecasts are an interesting starting point for future research, in particular also regarding the benefits of having an ensemble of NWP predictions available as input. Further, it would also be of interest to investigate whether post-processing methods could help to further improve the predictions of the IC-based UQ methods applied to deterministic data-driven weather forecasts we considered, for example by correcting some of the deficiencies observed for the GNP and IFSP approaches.
@@
Thus far, our comparisons have been focused on selected target variables, on using the gridded ERA5 data as ground truth, and on the CRPS as main evaluation metric. Operational weather services tend to evaluate their forecasts against the model’s own operational analysis, station observations (Rasp et al., 2024), and comprehensive comparisons of the UQ methods constitute an interesting starting point for future research. Over Europe, suitable station observation data has for example been collected within the EUPPBench benchmark dataset for post-processing (Demaeyer et al., 2023). Another important open question regarding the potential and limitations of data-driven weather models is whether they can reliably predict extreme weather events. Therefore, a targeted evaluation of the UQ methods in this regard, e.g., using proper weighted scoring rules (Lerch et al., 2017), represents another important direction for future model comparisons.
@@
The large data volumes and high dimensionality of global gridded predictions further poses a challenge regarding the scalability of ML-based post-processing methods such as DRN, for which it is an open question whether they will generalize well to global high-resolution forecasts. This calls for the development of new spatial post-processing methods which operate on the spatial forecast fields directly and are able to leverage predictive information present in the spatial structures, as well as for the development of suitable evaluation metrics. Over the past years, several approaches have been proposed, which utilize convolutional neural network or transformer architectures to enable probabilistic post-processing of spatial forecast fields (e.g., Grönquist et al., 2021; Ashkboos et al., 2022; Chapman et al., 2022; Ben Bouallègue et al., 2024b; Horat and Lerch, 2024). The recently introduced WeatherBench 2 dataset (Rasp et al., 2024) provides a useful framework for comparisons. In addition, future comparison should include inherently probabilistic data-driven models, such as GenCast (Price et al., 2023), NeuralGCM (Kochkov et al., 2023), or FuXi-ENS (Zhong et al., 2024). Although these methods already provide a data-driven ensemble forecast, their performance could potentially still be improved by applying additional post-processing for selected variables.